Government purchases (NOT including transfer payments) = $800
Transfer payments = $200
Total taxes = $1200
Output (income) = $5200
Consumption = $3000

Let's say the MPC = 0.75 and the government gives a tax cut of $20. As a result, private savings will                            [ Select ]                       ["fall", "not change", "rise"]         , public savings will                            [ Select ]                       ["fall", "rise", "not change"]         and this                            [ Select ]                       ["is not", "is"]         a good idea to stimulate economic investment.
Answer #1

Ans. Tax multiplier = -MPC/(1-MPC) = -0.75/(1-0.75) = -3

Change in output = tax multiplier * Change in tax

=> Change in output = -3*(-20)

=> Change in output = $60

a) Change in private savings = (1-MPC)*Change in output = (1-0.75)*60 = $15

Thus, savings increase as a result of tax cut

b) Change in public saving = Change in tax = -$20

Thus, public savings decrease by $20

c) No, it is not a good idea to stimulate economic investment because reduction in taxes increases the disposable income of households which increases the consumption spending which increases the transaction demand for money increasing interest rate in the economy which increases cost of borrowing and thus, reduces the investment spending in the economy.

* Please don’t forget to hit the thumbs up button, if you find the answer helpful.

Thank You